OpenStudy (happpily):

hiiiiii i need help. (Algebra 2)

OpenStudy (anonymous):

hi need help

OpenStudy (hysusonic):

whats the question

OpenStudy (anonymous):

0,8

OpenStudy (anonymous):

i don't want to help

OpenStudy (happpily):

oh ok thats fine...

OpenStudy (happpily):

0.8? um i really need help understanding this.. :o

OpenStudy (confluxepic):

I am taking Algebra 2 this year. I might be able to help in future questions.

OpenStudy (happpily):

oh okay awesome.thanks :3

OpenStudy (confluxepic):

You're welcome.

OpenStudy (anonymous):

medal

OpenStudy (happpily):

but you didn't help me......

OpenStudy (confluxepic):

I will try to help you. @Happpily @~

OpenStudy (happpily):

oh okay thank yooooooou!

OpenStudy (nikato):

Look at the three maximun coordinates given. To find the maximun value. Add the x and y coordinate for each pair and see which sum is the greatest.

OpenStudy (nikato):

Yu have (0,8) =0+8=8 (5,4)= 5+4=9 (9,0)= 9+0=9 So what's the maximun value?

OpenStudy (happpily):

(9, 0); 27 ? @nikato

OpenStudy (nikato):

No. It's just 9. Because 9 is the largest sum you get

OpenStudy (happpily):

oh okay thank you.

OpenStudy (nikato):

You're welcome